OpenStudy (anonymous):

Add and simplify. (2x2 – 11x) + (3x2 + 10x – 4) A. 5x4 – 4 B. 5x2 – x – 4 C. 6x2 – 11x – 4 D. 5x2 – 4

OpenStudy (hoblos):

(2x2 – 11x) + (3x2 + 10x – 4) = 2x2 – 11x + 3x2 + 10x – 4 add similar terms = (2+3)x2 +(-11+10)x – 4 = 5x2 – x – 4

OpenStudy (anonymous):

Thank You